OpenStudy (anonymous):

Solve: 5/3y - 1/2 = 5/6y -1/12

OpenStudy (anonymous):

10/6 y - 5/6 y = -1/12 + 6/12 5/6 y = 5/12 60 y= 30 y = 1/2